# DreamTalk: When Expressive Talking Head Generation Meets Diffusion Probabilistic Models

### v1

@@ B.2 More Results of Style Code Visualization
Figure 14: t-SNE visualization of style codes for 3 speakers, with darker hues representing increased emotional intensity.
We observe that each speaker’s style code distribution exhibits both common patterns and individualized characteristics. Common patterns include: Firstly, speaking styles of different emotions cluster together first, with styles of lower intensity being closer to neutral and those of higher intensity being further away. Secondly, speaking styles of anger and disgust, as well as fear and surprise, often cluster together, as shown in Fig. 14 (b) and (c). Note that unlike Ma et al. [46], our method does not incorporate losses to constrain style space.
Fig. 14 (a) illustrates an example of individualized characteristics. The speaker’s manifestation of fear closely resembles sadness, lacking the characteristic wide-eyed and open-mouthed expression, thereby positioning the speaking styles of fear nearer to those of sadness rather than surprise. Even within the same emotion, a speaker’s speaking style can exhibit notable variation. In Fig. 14 (c), the speaker’s dual expression of sadness—once with clenched teeth, similar to happy expressions, and another with depressed lip corners, akin to disgust—results in style codes close to the respective emotions. This observation diminishes the rationale for manually categorizing styles based on emotion and intensities in Ma et al. [46].
Appendix C Implementation Details
C.1 Architectual Details
Denoising Network.The audio encoder processes an input window of 11 sequential audio features, each of dimension $1024$. These features undergo dimension reduction to $256$ via a linear layer and then are fed to a transformer encoder comprising three 8-head transformer encoder layers, each with a hidden size of 256. Subsequently, a linear layer transforms the output tokens to yield audio tokens sized $11\times 256$. The audio tokens are concatenated with noisy motion and then added with the encoded diffusion step.
The style encoder ingests sequential expression parameters from style reference videos, each sequence sized $N\times 64$. These sequences, ranging in length from 64 to 256 frames, are initially expanded to 256 dimensions via a linear layer. Subsequently, they are introduced into a transformer encoder, composed of three 8-head layers, each with a hidden size of 256. The resulting output tokens, each with a dimension of $256$, are aggregated through self-attention pooling [56], yielding a style code of dimension $256$.
Within the decoder, the style code is repeated $11$ times, subsequently added with positional embedding to produce style tokens. These tokens, in conjunction with audio tokens, are processed by a transformer decoder, encompassing three 8-head layers, each with a hidden dimension of 256. Here, style tokens serve as the query, while audio tokens serve as both key and value. The middle output token is fed into a linear layer to predict facial motion.
Style-aware Lip Expert.
Figure 15: The architecture of mouth embedder and audio embedder. ConvBlock consists of a 1D CNN layer and a batch normalization layer. "k3, 32, s2" means that the kernel size is 3, the output dimension is 32, and the stride is 2. ConvRBlock is a ConvBlock with residual connection.
The face mesh is obtained by adding the mean shape to the product of the expression parameters and expression bases. Fig. 15 shows the architecture of the audio embedder and the mouth embedder.
Style Predictor. The style predictor is implemented as a transformer encoder comprising six 8-head transformer encoder layers, each with a hidden size of $256$. The input features are all linearly projected to $256$.
C.2 Data Details
C.2.1 Datasets
MEAD. The dataset, an in-lab talking-face corpus, features 60 speakers articulating eight emotions at three different intensity levels. When dividing the MEAD dataset into training and test subsets, we adhere to previously established methodologies [30].
HDTF. The dataset stands out as a high-resolution, in-the-wild audio-visual dataset. We designate $10\%$ of HDTF videos for testing and reserve the remainder for training.
Voxceleb2. Voxceleb2 is a large-scale talking head dataset with videos collected from YouTube. We redownload and recrop the videos to improve their resolution to $256\times 256$. Subsequently, approximately $80000$ high-quality videos are selected, with $400$ allocated for testing and the rest for training.
RAVDESS. The dataset features 24 professional actors (12 female, 12 male) vocalizing two lexically matched statements in a neutral North American accent. It encompasses a range of expressions in both speech and song, each articulated at two levels of emotional intensity, with an additional neutral expression included. We employ the speech data from RAVDESS, encompassing eight emotions, for evaluation.
C.2.2 Data Processing
The original videos are cropped and resized to $256\times 256$ pixels, aligning with the specifications in FOMM [59], and are sampled at 25 FPS. The 3DMM parameters are extracted by Deep3DFace [15].
Regarding the audio features used in the denoising network and the style-aware lip expert, we downsample the speech wave into the sampling rate of 16000 and extract acoustic features employing a pre-trained Wav2Vec2.0 model [2].
For audio features used in the style predictor, we extract them using a pre-trained HuBERT model [28]. Besides, we also utilize low-level audio features including Mel Frequency Cepstrum Coefficients (MFCC), Mel-filterbank energy features (FBANK), fundamental frequency, and voice flag. These two type of features are concatenated to represent the audio features used in the style predictor.
C.3 Training Details
Our framework is implemented on Pytorch. We employ Adam [32] for optimization, with a learning rate set to $0.0001$. The number of diffusion steps for the denoising network and style predictor is $1000$. The training batch size for the denoising network, style predictor, and lip expert is $64$, $64$, and $32$, respectively. $\lambda_{\text{denoise}}$, $\lambda_{\text{sync}}$, $n$, and $w$ are set to $1$, $1$, $5$, and $5$, respectively. The number of frames for style reference and audio used in style prediction is limited to $64-256$, corresponding to a time length of $2.56-10.24$ seconds. The denoising network, style predictor, and lip expert are trained on one NVIDIA Tesla A100 GPU for about $3$, $1$, and $10.5$ hours, respectively.
C.3.1 Finetuning PIRender
The renderer is fine-tuned with the losses in  Ren et al. [52] using MEAD. Instead of training with the self-reconstruction protocol where the source frame and target frame are from the same video, we select the source frame and target frame from the same speaker with different emotions. This enables the renderer to generate emotions different from the input portrait. This also allows our method to utilize portraits with emotions, unlike previous approaches that are confined to using neutral portraits [30]. We also observed that when fine-tuned only on the emotional talking head dataset, the renderer struggles with identity preservation. We argue that this problem stems from the fact that the number of speakers in the emotional dataset is limited. Therefore, we incorporate some neutral videos in Voxceleb into the data used for fine-tuning. This enhances the performance in identity preservation.
C.3.2 Training Style-aware Lip Expert
The style-aware lip expert is trained to discriminate whether the input audio and face motions are synchronized. We use cosine-similarity with binary cross-entropy loss to train the lip expert. Specifically, we compute cosine-similarity for the face motion embedding $\bm{e}^{m}$ and audio embedding $\bm{e}^{a}$ to represent the probability that the input audio-motion pair is synchronized. The training loss of the lip expert is:
$\mathcal{L}_{{\text{expert}}}={\text{BCE}}(\frac{\bm{e}^{m}\cdot\bm{e}^{a}}{\max(||\bm{e}^{m}||_{2}\cdot||\bm{e}^{a}||_{2},\epsilon)}),$ (10)
where $\epsilon$ is a small number for avoiding the division-by-zero error.
C.4 Inference Details
The inference of the denoising network can be accelerated with DDIM. We generate samples with $10$ DDIM steps and observe no performance drop. Generating a 30-second video offline takes $15.61$ seconds, with the face motion generation only taking $1.24$ seconds. During evaluation, the scale factor $\omega$ of classifier-free guidance is set to $1$. The style predictor uses the sampling algorithm of DDPM to predict style codes.
The emotion conveyed in the style reference (video or audio), should remain consistent to avoid confusing the model.
The head pose information, which is fed into the renderer, can be derived from real videos or generated using existing methods [83].
Appendix D Limitations and Future Work
@@
Despite DreamTalk’s promising advancements in expressive talking head generation, it encounters several challenges that open avenues for future research.
Firstly, the method occasionally produces artifacts, such as teeth flickering, around the mouth area, particularly during intense expressions. Generating teeth is a long-standing challenge in talking head generation since the algorithm needs to inpaint the teeth area that are often occluded in the input portrait. This problem is exacerbated under intense expressions where the teeth area expands. The issue can be mitigated by incorporating modules [44, 92, 36] proposed recently that enhance the teeth quality. A more comprehensive solution involves developing an emotion-specific renderer, as current renderers [90, 46, 20, 30] are mainly adaptations of existing facial reenactment methods with minimal modifications. An emotion-aware renderer would not only address the teeth generation issue but also enhance the overall expressiveness of emotions.
Secondly, DreamTalk does not account for the temporal variability in speaking styles. In real-life scenarios, a speaker’s style evolves over time, a feature our method currently overlooks. For instance, at the end of a speech, our method might still produce expressions of intense emotion, such as a wide-open mouth in surprise, instead of a more neutral, closed-mouth expression. Introducing a module that dynamically predicts speaking style over time could address this limitation.
@@
Thirdly, the style predictor sometimes struggles with accurately identifying emotions in low-intensity audio clips from the MEAD dataset (some audio in MEAD intensity level 1 videos). The emotion conveyed in these audio clips is very similar to neutral emotion and thus confusing the prediction. We also observed that although the videos in the MEAD dataset clearly express emotions due to the good training and supervision of the speakers, sometimes the audio does not correspond with the emotions that should be expressed. Therefore, To enhance prediction accuracy, employing a dataset where the audio closely aligns with the expressed emotions could be beneficial. Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [85, 8].
@@
Despite these challenges, DreamTalk marks a significant stride in the realm of high-quality, expressive talking head generation, setting a foundation for further innovations.
Appendix E Ethical Consideration
DreamTalk is able to generate realistic talking head videos. This positions DreamTalk with a broad spectrum of potential applications, each carrying intricate societal implications. While DreamTalk holds significant potential in amplifying and enriching human creative endeavors and may pave the way for innovative tools for creative professionals, its capabilities also harbor risks. There’s a possibility for DreamTalk to generate content that might encompass or imply sexual themes, promote hatred, or depict violence. Misuse of DreamTalk could lead to negative repercussions on individuals or groups, potentially erasing or maligning them, perpetuating stereotypes, and subjecting them to disrespect. Other concerns include the potential for harassment, intimidation, or exploitation. Furthermore, DreamTalk’s capabilities might be harnessed to mislead or spread misinformation.
Before releasing DreamTalk, we have implemented and plan to introduce several safeguards to curb potential misuse. We’ve purged detrimental content from the training dataset and incorporated visual filters to deter users from creating harmful outputs. To counteract biases in the generated results, we’re enhancing the dataset’s diversity by manually ensuring balance, which will reduce instances of erasure, stereotype perpetuation, indignity, and uneven quality across inputs. Users will be advised against using images without the depicted individuals’ consent to combat harassment and bullying. To prevent the spread of misinformation, all DreamTalk outputs will bear watermarks indicating their synthetic nature. Our pre-release strategy involves a thorough risk assessment, leveraging a growing suite of safety evaluations and red teaming techniques. We’ll also scrutinize the findings from pilot tests centered on new use cases and conduct in-depth post-release assessments. Both automated and manual monitoring mechanisms are in development to preempt misuse. Our commitment remains steadfast in continuously researching ways to minimize adverse societal effects.

### v2

@@ IV-D Style Code Visualization
Figure 13: t-SNE visualization of style codes. Left: Style codes from 15 speakers. Each color indicates style codes from an identical speaker. Right: Style codes from a speaker, with darker hues representing increased emotional intensity.
Figure 14: Modulating style intensity by adjusting the scale $\omega$ of classifier-free guidance.
We use t-SNE [103] to map style codes from the MEAD dataset’s 15 speakers into a 2D space. Each speaker exhibits 22 distinct speaking styles, comprising seven emotions at three intensity levels, alongside a neutral style. For each style of each speaker, we randomly select 10 videos to extract style codes for visualization.
The left in fig. 13 shows that style codes first cluster based on identity rather than emotion, indicating the correlation between speaker identity and speaking styles, hence justifying the rationale for using portrait information to infer speaking styles.
The right in fig. 13 shows that even within the same emotion and the same intensity, a speaker’s expressions can vary significantly. The speaker expresses intense sadness in two ways. In one way, the speaker clenches teeth (top left portrait), similar to happy expressions (bottom left portrait), while in another way, the speaker depresses lip corners (bottom right portrait). The observation suggests StyleTalk’s assumption that speaking styles are consistent in videos with the same emotions is incorrect, which may impair performance.
IV-E Modulating Style Intensity
As elaborated in section III-C, the scale factor $\omega$ in the classifier-free guidance scheme can modulate the intensity of any input style, even those unseen during training. As shown in fig. 14, the style reference is in-the-wild and adjusting $\omega$ either amplifies or attenuates the designated style. When $\omega=0$, DreamTalk produces a talking head with a neutral expression. We observed a noticeable decline in lip-sync accuracy when the scale factor $\omega$ exceeds $2$.
IV-F User Study
TABLE VI: Mean ratings of user study with $95\%$ confidence intervals.
[…]
MakeItTalk [55] 2.73 $\pm$ 0.08 3.10 $\pm$ 0.08 -
Wav2Lip [54] 3.46 $\pm$ 0.07 2.45 $\pm$ 0.05 -
PC-AVS [59] 3.29 $\pm$ 0.04 2.86 $\pm$ 0.07 -
AVCT [47] 3.43 $\pm$ 0.07 3.39 $\pm$ 0.08 -
GC-AVT [3] 3.40 $\pm$ 0.06 2.21 $\pm$ 0.05 2.70 $\pm$ 0.05
EAMM [5] 2.83 $\pm$ 0.08 2.46 $\pm$ 0.07 2.43 $\pm$ 0.06
StyleTalk [2] 3.42 $\pm$ 0.05 3.48 $\pm$ 0.07 2.88 $\pm$ 0.07
SadTalker [60] 3.44 $\pm$ 0.07 3.51 $\pm$ 0.06 -
PD-FGC [1] 3.16 $\pm$ 0.06 2.80 $\pm$ 0.06 3.06 $\pm$ 0.07
EAT [6] 3.39 $\pm$ 0.07 2.87 $\pm$ 0.08 2.91 $\pm$ 0.07
DreamTalk 3.80 $\pm$ 0.07 3.77 $\pm$ 0.07 3.34 $\pm$ 0.08
Ground Truth 4.21 $\pm$ 0.08 4.09 $\pm$ 0.08 -
Emotional Talking Head Generation. We conduct a user study to further evaluate our method. We generate 30 test samples for each method, which cover diverse speaking styles and speakers, and recruit 22 participants to rate samples. Each participant is required to rate all samples (from 1 to 5, 5 is the best) on three aspects: (1) lip sync quality, (2) video realness, and (3) style consistency between the generated videos and the style reference (This metric is only evaluated on emotional methods. Since Ground Truth videos do not express the expressions reflected in the style references. The score for Ground Truth is omitted.). As shown in table VI, our method outperforms baselines across all aspects. A one-way ANOVA and a post-hoc Tukey test identify a significant difference ($p<0.001$) between our method and other baselines on all aspects.
Speaking Style Prediction. In our user study, we evaluate the alignment between the original and predicted speaking styles. Directly assessing the alignment of speaking styles can be somewhat ambiguous, so we employ a comparative approach for evaluation. Specifically, we create a series of video triplets. Each triplet consisted of a test video from our dataset and two generated videos. The first video was generated using a style code predicted from an input portrait, sharing the same speaker identity as in the test video but displaying a neutral emotion, combined with the audio from the test video. The second video is generated using the style code extracted from videos with the same emotion but from a speaker different from the one in the test video. We recruit 20 participants. Each participant is then asked to evaluate 20 triplets and identify which of the generated videos most accurately reflected the speaking style of the test video. The videos generated using predicted style codes are preferred in $75.8\%$ of all ratings. This indicates that the style predictor is able to infer personalized speaking styles that are aligned with the audio.
V Limitations
@@
Despite DreamTalk’s promising advancements in emotional talking head generation, it encounters several challenges that open avenues for future research.
When using a constant style reference, DreamTalk generates expressions that are strictly consistent with the main expressions in the reference but lack expression changes over time, such as eye blinking. Generating emotions with rich temporal variations is, compared to methods that generate neutral or coarse-grained emotions, more difficult for methods that achieve fine-grained emotional control, like DreamTalk. This is because these methods must achieve both precise control of expressions and diverse changes. To address the issue, temporal changes in expressions can be achieved by using temporally changing style references when generating each video frame. As discussed in Supp. Mat., DreamTalk can achieve smooth expression changes by changing the style references. Eye blinks can be achieved by using eye blink loss [60] during training or post-editing 3DMM, a common practice used in previous methods [104, 66] that also aim to control expressions. Specifically, we can obtain the parameter changes of blinking and then edit the generated expression parameters. A video with the eye blinking through post-editing is shown in Supp. Video.
DreamTalk may change the speaker’s identity when the identity in the reference video highly differs from that in the portrait. The reason is that 3DMM expression parameters leak identity information. This leakage leads to the generated identity becoming somewhat similar to the identity in the reference. The issue can be mitigated by adopting 3DMM which decouples expression and identity information more effectively.
@@
The style predictor sometimes struggles with accurately identifying emotions in low-emotion-intensity audio clips from the MEAD dataset. The reason is that in some MEAD videos, the audio does not correspond with the expressed emotions. To enhance prediction accuracy, it is beneficial to employ a dataset where the audio closely aligns with the expressed emotions. Another solution is to incorporate textual information from audio during prediction, a strategy commonly employed in speech emotion recognition [105, 106].
DreamTalk occasionally produces artifacts around the mouth area, such as teeth flickering, particularly during intense expressions. The issue comes from the renderer and can be mitigated by using more advanced renderers.
@@
Despite these challenges, DreamTalk marks a significant stride in the realm of emotional talking head generation, paving the way for further innovations.
VI Conclusion
In this work, we propose DreamTalk, a novel diffusion-based framework that can consistently generate high-quality talking heads in diverse speaking styles and conveniently use audio to specify personalized emotions. We develop a denoising network for creating emotional, audio-driven facial motions and introduce a style-aware lip expert to optimize lip-sync while maintaining emotion intensity. Additionally, we devise a style predictor that infers speaking styles directly from audio, eliminating the need for video references. Extensive experiments validate the efficacy of DreamTalk. The results demonstrate that employing diffusion models markedly improves the quality of emotional talking head generation.
VII Ethical Consideration
DreamTalk can generate vivid talking head videos, opening up diverse applications but also posing risks like promoting hatred or depicting violence. Misuse could harm individuals or groups, perpetuate stereotypes, and spread misinformation. To mitigate these risks, we’ve implemented safeguards like watermarks on all outputs and advising against using images without consent. We remain committed to continuous research to minimize adverse societal impacts.